OpenStudy (anonymous):

a student opens a mathematics book to two facing pages. The product of the page numbers is 506. Find the page numbers I got 22/23

OpenStudy (anonymous):

using the quadratic equation i got x=44/2 and x=-46/2

OpenStudy (zarkon):

22 and 23 are correct

OpenStudy (apoorvk):

yeah 22 and 23 are correct

OpenStudy (anonymous):

cool. I really hate these quadratic equations lol. Been doing them all day and am so burned out on them lol